crizzyyy1063:

A carpet installer charges a flat rate of $40, plus $1.50 per square foot for labor. So, the total cost for labor depends on the amount of carpet installed. someone do a zoom plzzz

glizxzy:

x > 0 domain y = 40+1.5x relationship y > 40 range the greater than signs have the line under them i just dont know how to put that there

dle28022:

Domain y = 40+1.5x relationship range explanaton: The domain is the amount of carpet installed. We can install zero carpeting or any positive amount of carpeting. Now writing the equation for the cost They charge 40 flat rate plus 1.50 per square foot y = 40+1.5x Let x = 0 y = 40+0 Even if they do not install anything, they still charge 40 dollars flat rate. That is the minimum amount we have to pay. As x increases, the cost will increase. That is the range, or the values of y